evaluate: sin200*cos70+cos200*sin70
a circle with a radius of 6 is inscribed in a square. if a dart is thrrown what is the probability that it lands inside the circle?
state the domain that is needed for tanx so that its inverse Arctanx is a function. determine the range of Arctanx

Answer by user_dude2008(1862)

sin200*cos70+cos200*sin70 = sin(200+70) = sin(270) = -1